OpenStudy (anonymous):

What is the equation of the line, in general form, that passes through the point (-1, -1) and is parallel to the line whose equation is x + y = 3?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=x+%2B+y+%3D+3

OpenStudy (anonymous):

that might help

OpenStudy (oaktree):

Rather than giving a website, let's try to work it out together. First we need to figure out the slope of our line, and since we have a different line it's parallel to, we can use that to figure out our slope. x+y=3 in point-slope form is y=-x+3, so our slope is -1. Now to figure out our y-intercept, we plug in the point given into the point-slope equation and we gave -1=-1(-1)+b, which gives us b=-2. So the equation is y=-x-2. Makes sense?

OpenStudy (anonymous):

x - y + 2 = 0 x + y + 2 = 0 x + y - 2 = 0 These are the options I have to choose from...

OpenStudy (oaktree):

So it's x+y+2=0.

OpenStudy (anonymous):

Oh ok thank you so much

OpenStudy (oaktree):

No problem.